A Juggler tosses a ball up in the air initial speed u. At the instant it reaches its maximum height H, he tosses up a second ball with the same initial speed. In what height will two ball collide?

Grade:12

1 Answers

Deepak Sangle
6 years ago
For first ball,0 = u^2 -2gH thus u^2 = 2gH ...... (1)after the ball collides at height h, for first ball H-h = 1/2 gt^2 .....(2) and for second,h =ut - 1/2 gt^2 ......(3) from eqn (2) and (3) we get H = ut or H^2 = u^2 t^2 or H^2 = 2gHt^2 .... (from 1) and thus gt^2 = H /2 Putting it in eqn (3) we get h = H - 1/2 H/2 and thus h = 3H / 4